OpenStudy (anonymous):

the sides of a rectangle is the square root of 14 and the square root of 20.. what is the area? and if it can be, please put it in radical form.

OpenStudy (gina):

\[L timesB =AREA,SO \sqrt{14}\times \sqrt{20}=280^{1/4}\]

OpenStudy (gina):

THIS THE answere

OpenStudy (anonymous):

what would that be in radical form?

OpenStudy (gina):

16.73320053

OpenStudy (gina):

what are the units?

OpenStudy (gina):

16.73320053,,,this the anwer in radical